# Computational Caustic Design for Surface Light Source

### v1

@@ 1 Introduction
The interaction between light and objects can produce various captivating illumination effects, such as caustic phenomena that occur when light undergoes refraction or reflection. Precisely controlling this caustic process by designing freeform surfaces has important applications in computer graphics [1, 2] and geometric optics [3, 4]. Additionally, accurately controlling the caustic process to achieve desired flux distributions has broad applications in fields such as art design, interior lighting, and industrial manufacturing.
In the field of non-imaging optics, the shape of freeform surface lenses is typically designed to control various properties of light in order to create the desired illumination effects. Freeform surfaces here refer to optical surfaces without linear or rotational symmetry [5]. They offer a high degree of freedom, thereby enabling precise control over light beams.
Existing methods in computer graphics typically reconstruct the final freeform surface by calculating certain shape characteristics of the required lens, such as its normals [1] or visibility diagram [2]. However, these methods almost always assume Lambertian point light sources or parallel light [6, 1, 2], which are inaccurate for real-world scenarios as ideal point or parallel light sources do not exist. Instead, real light sources often have a finite area. This discrepancy affects the accuracy of caustic design, causing the designed lenses to produce blurred images on the receiving plane.
@@
Additionally, existing methods for simulating surface light source illumination typically sample rays based on the actual light intensity distribution on the surface [7, 8, 9]. This deviates from true physical behavior and incurs high computational costs. More critically, the exact light intensity distribution of an LED surface light source is often unknown, making it impossible for ray-sampling-based methods to accurately simulate illumination [10].
@@
It is worth noting that optimizing the parameters of the surface light source model with the help of caustic lenses with known shapes is a crucial aspect of our design. In the optimization process of the model, the introduction of caustic lenses can amplify the differences in the propagation directions of various light rays, clearly highlighting the illumination characteristics of the surface light source. At the same time, using caustic lenses can form meaningful patterns on the receiving plane, making it easier for us to evaluate image details. To address the issue of imprecision in ray sampling strategies, we introduced the concept of light flux. By calculating the light flux within a region, we achieve an exact simulation of the light propagation process. This light energy transfer approach is more accurate in a physical sense and does not require prior knowledge of the intensity distribution on the light source surface when combined with our light source model. Furthermore, considering that the inherent size of the surface light source introduces boundary constraints, we designed a contraction mapping [11] to transform the constrained optimization problem into an unconstrained one, which can be efficiently solved by using the L-BFGS algorithm [12]. To further reduce the time cost of the rendering process, we incorporated a parallel design into the optimization process.
This paper addresses a fundamental limitation in computational caustic design: the rendering blur caused by idealized point or parallel light source assumptions. By introducing a physically-based, optimizable model for surface light sources, our method significantly improves the accuracy of caustic generation (see Figure 1). This advance is critical for a variety of practical applications where precision is important, including artistic design [13], industrial lighting [14, 15], solar energy collection [16, 17, 18], and the design of advanced optical instruments and lenses for medical, biochemical [19, 20], and astronomical applications [21].
@@ 2 Related Work
Freeform caustic surface design.  Freeform surfaces are special optical devices commonly used to achieve beam control. Specifically, by designing a high degree of freedom optical surface (usually a lens), the light emitted by the light source is redirected to produce a specified irradiance distribution [22]. Compared with traditional optical surfaces (including spherical, non-spherical, and cylindrical surfaces), due to the huge degree of freedom of freeform surfaces, the performance of optical systems can reach new heights in fields such as imaging [23, 24, 25], illumination [26, 27], and laser beam shaping [28]. In fact, many real-life scenarios require specialized optical systems, such as automobile headlights [14] and projectors [29, 30]. Thus, it can be seen that freeform surfaces play an important role in achieving specific control of light beams [31, 32].
@@
Another type of work was first proposed by Finckh et al. [33]. They used freeform surfaces to refract or reflect light to form the desired caustic image. This work uses a stochastic approximation optimization algorithm to determine the surface geometry from the caustics. This requires a large amount of computation and cannot achieve accurate control of areas with weak light intensity. Subsequent work starts from the principle of local energy conservation and formulates it as a problem of solving a class of Monge-Ampère equations [14, 34, 35]. These methods assume that the lens shape is a continuous and smooth surface, thereby reducing artifacts to achieve higher-quality images. Some other works solve the shape of the freeform surface by means of the integral normal field [13] or formulate it as a Poisson reconstruction problem [36]. However, the freeform surfaces obtained based on these methods cannot produce high-contrast images. Schwartzburg et al. [1] solved this problem. They obtained the lens surface by calculating the normal of the lens surface through optimal transport, which can produce a pure black area on the receiving plane. Meyron et al. [2] proposed an algorithm to solve the optimal transport problem using visibility maps, further reducing the loss of light energy. Marc et al. [37] combined caustic design with novel neural denoising components and a learned gradient descent optimizer, improving the stability and quality of the iterative reconstruction process.
@@
Differentiable rendering.  Differentiable rendering is included in the field of inverse rendering, which estimates physical parameters from imaging data using a differentiable renderer. These rendering methods describe the image formation process in a differentiable way and can effectively calculate the gradients of relevant free parameters. Therefore, they can be used in parameter search strategies based on gradient descent. In caustic design, differentiable rendering is used to obtain images generated by the joint action of illumination sources and free lenses, as well as the gradients of relevant physical parameters. However, traditional rendering methods usually include non-differentiable components. Rasterization [38] is a typical example. Therefore, it is not easy to directly optimize the physical properties in the rendering scene based on the simulation results. In recent years, many works have been studying how to make the rendering process differentiable. Liu et al. [39] proposed soft rasterization, which uses probability distributions and aggregation functions to make the rendering pipeline differentiable. Other works [40, 41, 42, 43] solve the problem of computing gradients with respect to the scene geometry, which is not differentiable due to the discontinuity of the mesh. After that, some differentiable renderers have emerged, such as Mitsuba2 [44, 45, 46], Mitsuba3 [47, 48], and Path Replay Backpropagation [49]. The ray tracing-based approach in Mitsuba3 can be applied to caustic design. However, since this renderer is not specifically tailored for caustic design, there remains considerable room for improvement in its rendering results. Wang et al. [50] proposed a strategy based on ray tracing and deep lens systems, which can be applied to a variety of inverse design problems. However, certain limitations still remain regarding its performance and applicability.
@@
In addition to path tracing algorithms, photon mapping is also a classic global illumination algorithm. Originally proposed by Jensen in 1996 [51], it efficiently simulates complex lighting effects such as caustics by tracing photon paths and constructing a photon map. To optimize memory usage, Hachisuka et al. [52] introduced Progressive Photon Mapping (PPM), which performs photon tracing in multiple iterations. Several other studies have also introduced various improvements to the photon mapping algorithm. Stochastic Progressive Photon Mapping (SPPM) [53] enabled the support of effects such as glossy reflections. Georgiev et al. [54] and Hachisuka et al. [55] reformulated photon mapping as a vertex merging technique compatible with path integration, which laid the foundation for differentiable photon mapping. Xing et al. [56] leveraged Extended Path Space Manifolds (EPSMs) to achieve differentiability of path vertex positions and color contributions with respect to the scene parameters. However, the rendered results still exhibit significant noise and continue to require light path sampling. To achieve a physically more accurate caustic design, we use a differentiable framework based on flux transport.
@@
Optimal Transport in Caustic Design.  The problem of caustic lens design is essentially an issue of adjusting the light field distribution. Finding the correspondence between two density distributions is an important problem in mathematics, and optimal transport (OT) is an effective tool for solving this problem. It finds an optimal way to transfer one mass distribution into another, minimizing the total transportation cost. This problem was first proposed by the French mathematician Gaspard Monge in 1781 [57] and was later extended by Kantorovich to allow for the splitting of mass [58]. The classic optimal transport problem includes both discrete and continuous forms. Subsequently, an important semi-discrete form emerged, which seeks an optimal mapping that converts a discrete distribution into a continuous one [59, 60]. [1, 61, 62, 63] all adopt the optimal transport strategy. However, this strategy is only suitable for caustics based on point light sources or parallel light. Taking the triangle mesh representation of a lens as an example, when the input light source is a surface light source, each triangular area on the lens surface no longer controls only a single light ray. Our approach solves the above issue.
@@ 3.4 Optimization design for light source model
Contraction function.  In the above optimization problem (18), the penalty for boundary constraints does not necessarily ensure that these parameters are strictly controlled within the boundaries. Furthermore, the selection of the coefficient of each penalty term has a significant impact on the optimization process and may lead to early termination. Inspired by [11], we design a monotonic and continuously differentiable contraction function for the position parameters:
@@
To solve unconstrained optimization problems (21), we use the L-BFGS algorithm [12] to compute the numerical solution and leverage GPU parallel computation for the target value and its gradient (discussed in the next section). Due to the scale and complexity of the aforementioned optimization problem, we use an automatic differentiation data structure to precisely evaluate the gradient after each iteration.
@@ 3.5 Caustic design based on optimized model
After obtaining the optimized light source model, our current goal is to optimize the lens shape so that the caustic images formed under the combination of these point light sources are closely aligned with the given target images (see Figure 11). Given the high complexity and nonlinearity of this optimization, it would be challenging to optimize the lens shape directly from a planar surface. We comprehensively consider flux consistency and normal integrability to obtain the initial shape of the lens. Then, for optimization, we adapt the design of several loss terms from [64].
@@
where $H_{i}$ is the discrete mean curvature for the face $t_{i}$, $A_{i}$ is the area of $t_{i}$. Specific details can be referred to [64].
@@ 4.4 Caustic design
Our analysis ignores errors from reflection, absorption, and surface imperfections, an assumption justified by our use of acrylic (PMMA) for the physical prototypes. This material was chosen for its ideal optical and mechanical properties. In the visible spectrum, acrylic exhibits very low reflectance and near-zero absorption, which minimizes energy loss during light transport. Furthermore, its excellent wear resistance and machinability allow for a high-fidelity surface finish that reduces the impact of fabrication-related imperfections [65, 66]. Since these secondary physical effects are demonstrably minimal in our experimental setup, it is a valid and well-justified simplification for our model to focus on controlling light transport purely through the refractive properties of the lens geometry.
@@ 4.5 Effectiveness of algorithm components
We conducted comparative experiments with Mitsuba3 [47]. Mitsuba3 uses a path tracing algorithm to simulate the light transport process, which is a variant of ray tracing. It employs a Monte Carlo estimator to compute integral values by randomly emitting multiple rays to evaluate light intensity, thereby simulating the propagation of light in space. This method requires sampling light rays, which has certain limitations in computational efficiency. Moreover, Mitsuba3 is not specifically designed for caustic design, and therefore its performance in terms of contrast and accuracy is inferior to our method (see Figure 13). It also completely neglects the consideration of actual fabrication in the overall design process.

### v2

@@ 1 Introduction
The interaction between light and objects can produce various captivating illumination effects, such as caustic phenomena that occur when light undergoes refraction or reflection. Precisely controlling this caustic process by designing freeform surfaces has important applications in computer graphics [schwartzburg2014high, meyron2018light] and geometric optics [budhu2019novel, kass2021]. Additionally, accurately controlling the caustic process to achieve desired flux distributions has broad applications in fields such as art design, interior lighting, and industrial manufacturing.
In the field of non-imaging optics, the shape of freeform surface lenses is typically designed to control various properties of light in order to create the desired illumination effects. Freeform surfaces here refer to optical surfaces without linear or rotational symmetry [reimers2017freeform]. They offer a high degree of freedom, thereby enabling precise control over light beams.
Existing methods in computer graphics typically reconstruct the final freeform surface by calculating certain shape characteristics of the required lens, such as its normals [schwartzburg2014high] or visibility diagram [meyron2018light]. However, these methods almost always assume Lambertian point light sources or parallel light [Shen21, schwartzburg2014high, meyron2018light], which are inaccurate for real-world scenarios as ideal point or parallel light sources do not exist. Instead, real light sources often have a finite area. This discrepancy affects the accuracy of caustic design, causing the designed lenses to produce blurred images on the receiving plane.
@@
Additionally, existing methods for simulating surface light source illumination typically sample rays based on the actual light intensity distribution on the surface [xu2025hybrid, zli2020, sun2021end]. This deviates from true physical behavior and incurs high computational costs. More critically, the exact light intensity distribution of an LED surface light source is often unknown, making it impossible for ray-sampling-based methods to accurately simulate illumination [raytracing1984].
@@
It is worth noting that optimizing the parameters of the surface light source model with the help of caustic lenses with known shapes is a crucial aspect of our design. In the optimization process of the model, the introduction of caustic lenses can amplify the differences in the propagation directions of various light rays, clearly highlighting the illumination characteristics of the surface light source. At the same time, using caustic lenses can form meaningful patterns on the receiving plane, making it easier for us to evaluate image details. To address the issue of imprecision in ray sampling strategies, we introduced the concept of light flux. By calculating the light flux within a region, we achieve an exact simulation of the light propagation process. This light energy transfer approach is more accurate in a physical sense and does not require prior knowledge of the intensity distribution on the light source surface when combined with our light source model. Furthermore, considering that the inherent size of the surface light source introduces boundary constraints, we designed a contraction mapping [barron2022mipnerf360] to transform the constrained optimization problem into an unconstrained one, which can be efficiently solved by using the L-BFGS algorithm [Liu1989OnTL]. To further reduce the time cost of the rendering process, we incorporated a parallel design into the optimization process.
This paper addresses a fundamental limitation in computational caustic design: the rendering blur caused by idealized point or parallel light source assumptions. By introducing a physically-based, optimizable model for surface light sources, our method significantly improves the accuracy of caustic generation (see Figure 1). This advance is critical for a variety of practical applications where precision is important, including artistic design [kiser2013architectural], industrial lighting [zhu2013optical, kim2013design], solar energy collection [wallhead2012design, xie2011concentrated, leutz2012nonimaging], and the design of advanced optical instruments and lenses for medical, biochemical [goure2013optics, rovati1996design], and astronomical applications [bland2009astrophotonics].
@@ 2 Related Work
Freeform caustic surface design.  Freeform surfaces are special optical devices commonly used to achieve beam control. Specifically, by designing a high degree of freedom optical surface (usually a lens), the light emitted by the light source is redirected to produce a specified irradiance distribution [winston2005nonimaging]. Compared with traditional optical surfaces (including spherical, non-spherical, and cylindrical surfaces), due to the huge degree of freedom of freeform surfaces, the performance of optical systems can reach new heights in fields such as imaging [Rolland:21, Mao:18, Cheng:21], illumination [Rengmao2018Design, Li18], and laser beam shaping [Feng2021Iterative]. In fact, many real-life scenarios require specialized optical systems, such as automobile headlights [zhu2013optical] and projectors [damberg2016high, zhao2013integral]. Thus, it can be seen that freeform surfaces play an important role in achieving specific control of light beams [Duerr15, Li18].
@@
Another type of work was first proposed by Finckh et al. [finckh2010geometry]. They used freeform surfaces to refract or reflect light to form the desired caustic image. This work uses a stochastic approximation optimization algorithm to determine the surface geometry from the caustics. This requires a large amount of computation and cannot achieve accurate control of areas with weak light intensity. Subsequent work starts from the principle of local energy conservation and formulates it as a problem of solving a class of Monge-Ampère equations [zhu2013optical, ZHANG2014297, Chang_2016]. These methods assume that the lens shape is a continuous and smooth surface, thereby reducing artifacts to achieve higher-quality images. Some other works solve the shape of the freeform surface by means of the integral normal field [kiser2013architectural] or formulate it as a Poisson reconstruction problem [yue2014poisson]. However, the freeform surfaces obtained based on these methods cannot produce high-contrast images. Schwartzburg et al. [schwartzburg2014high] solved this problem. They obtained the lens surface by calculating the normal of the lens surface through optimal transport, which can produce a pure black area on the receiving plane. Meyron et al. [meyron2018light] proposed an algorithm to solve the optimal transport problem using visibility maps, further reducing the loss of light energy. Marc et al. [Kassubeck2021NSfCRA] combined caustic design with novel neural denoising components and a learned gradient descent optimizer, improving the stability and quality of the iterative reconstruction process.
@@
Differentiable rendering.  Differentiable rendering is included in the field of inverse rendering, which estimates physical parameters from imaging data using a differentiable renderer. These rendering methods describe the image formation process in a differentiable way and can effectively calculate the gradients of relevant free parameters. Therefore, they can be used in parameter search strategies based on gradient descent. In caustic design, differentiable rendering is used to obtain images generated by the joint action of illumination sources and free lenses, as well as the gradients of relevant physical parameters. However, traditional rendering methods usually include non-differentiable components. Rasterization [pineda1988parallel] is a typical example. Therefore, it is not easy to directly optimize the physical properties in the rendering scene based on the simulation results. In recent years, many works have been studying how to make the rendering process differentiable. Liu et al. [liu2019soft] proposed soft rasterization, which uses probability distributions and aggregation functions to make the rendering pipeline differentiable. Other works [Bangaru2020, Li2012Diff, Zhang2020Path, Zhang2021PSDR] solve the problem of computing gradients with respect to the scene geometry, which is not differentiable due to the discontinuity of the mesh. After that, some differentiable renderers have emerged, such as Mitsuba2 [Loubet2019, NimierDavid2020, NimierDavidVicini2019], Mitsuba3 [Mitsuba3, Jakob2020DrJit], and Path Replay Backpropagation [Vicini2021]. The ray tracing-based approach in Mitsuba3 can be applied to caustic design. However, since this renderer is not specifically tailored for caustic design, there remains considerable room for improvement in its rendering results. Wang et al. [wang2022differentiable] proposed a strategy based on ray tracing and deep lens systems, which can be applied to a variety of inverse design problems. However, certain limitations still remain regarding its performance and applicability.
@@
In addition to path tracing algorithms, photon mapping is also a classic global illumination algorithm. Originally proposed by Jensen in 1996 [jensen1996global], it efficiently simulates complex lighting effects such as caustics by tracing photon paths and constructing a photon map. To optimize memory usage, Hachisuka et al. [Hachisuka2008] introduced Progressive Photon Mapping (PPM), which performs photon tracing in multiple iterations. Several other studies have also introduced various improvements to the photon mapping algorithm. Stochastic Progressive Photon Mapping (SPPM) [Hachisuka2009] enabled the support of effects such as glossy reflections. Georgiev et al. [Georgiev2012] and Hachisuka et al. [Hachisuka2012] reformulated photon mapping as a vertex merging technique compatible with path integration, which laid the foundation for differentiable photon mapping. Xing et al. [xing2024differentiable] leveraged Extended Path Space Manifolds (EPSMs) to achieve differentiability of path vertex positions and color contributions with respect to the scene parameters. However, the rendered results still exhibit significant noise and continue to require light path sampling. To achieve a physically more accurate caustic design, we use a differentiable framework based on flux transport.
@@
Optimal Transport in Caustic Design.  The problem of caustic lens design is essentially an issue of adjusting the light field distribution. Finding the correspondence between two density distributions is an important problem in mathematics, and optimal transport (OT) is an effective tool for solving this problem. It finds an optimal way to transfer one mass distribution into another, minimizing the total transportation cost. This problem was first proposed by the French mathematician Gaspard Monge in 1781 [monge1781memoire] and was later extended by Kantorovich to allow for the splitting of mass [Kantorovich2006Translocation]. The classic optimal transport problem includes both discrete and continuous forms. Subsequently, an important semi-discrete form emerged, which seeks an optimal mapping that converts a discrete distribution into a continuous one [levy2015numerical, merigot2011multiscale]. [schwartzburg2014high, Oliker:07, Oliker2011Designing, Feng2021Iterative] all adopt the optimal transport strategy. However, this strategy is only suitable for caustics based on point light sources or parallel light. Taking the triangle mesh representation of a lens as an example, when the input light source is a surface light source, each triangular area on the lens surface no longer controls only a single light ray. Our approach solves the above issue.
@@ 3.4 Optimization design for light source model
Contraction function.  In the above optimization problem (18), the penalty for boundary constraints does not necessarily ensure that these parameters are strictly controlled within the boundaries. Furthermore, the selection of the coefficient of each penalty term has a significant impact on the optimization process and may lead to early termination. Inspired by [barron2022mipnerf360], we design a monotonic and continuously differentiable contraction function for the position parameters:
@@
To solve unconstrained optimization problems (21), we use the L-BFGS algorithm [Liu1989OnTL] to compute the numerical solution and leverage GPU parallel computation for the target value and its gradient (discussed in the next section). Due to the scale and complexity of the aforementioned optimization problem, we use an automatic differentiation data structure to precisely evaluate the gradient after each iteration.
@@ 3.5 Caustic design based on optimized model
After obtaining the optimized light source model, our current goal is to optimize the lens shape so that the caustic images formed under the combination of these point light sources are closely aligned with the given target images (see Figure 11). Given the high complexity and nonlinearity of this optimization, it would be challenging to optimize the lens shape directly from a planar surface. We comprehensively consider flux consistency and normal integrability to obtain the initial shape of the lens. Then, for optimization, we adapt the design of several loss terms from [sun2025end].
@@
where $H_{i}$ is the discrete mean curvature for the face $t_{i}$, $A_{i}$ is the area of $t_{i}$. Specific details can be referred to [sun2025end].
@@ 4.4 Caustic design
Our analysis ignores errors from reflection, absorption, and surface imperfections, an assumption justified by our use of acrylic (PMMA) for the physical prototypes. This material was chosen for its ideal optical and mechanical properties. In the visible spectrum, acrylic exhibits very low reflectance and near-zero absorption, which minimizes energy loss during light transport. Furthermore, its excellent wear resistance and machinability allow for a high-fidelity surface finish that reduces the impact of fabrication-related imperfections [PMMA2016review, PMMA2014Study]. Since these secondary physical effects are demonstrably minimal in our experimental setup, it is a valid and well-justified simplification for our model to focus on controlling light transport purely through the refractive properties of the lens geometry.
@@ 4.5 Effectiveness of algorithm components
We conducted comparative experiments with Mitsuba3 [Mitsuba3]. Mitsuba3 uses a path tracing algorithm to simulate the light transport process, which is a variant of ray tracing. It employs a Monte Carlo estimator to compute integral values by randomly emitting multiple rays to evaluate light intensity, thereby simulating the propagation of light in space. This method requires sampling light rays, which has certain limitations in computational efficiency. Moreover, Mitsuba3 is not specifically designed for caustic design, and therefore its performance in terms of contrast and accuracy is inferior to our method (see Figure 13). It also completely neglects the consideration of actual fabrication in the overall design process.
@@ Acknowledgments
[Uncaptioned image] Sizhuo Zhou is currently working toward a PhD degree in the School of Mathematical Sciences, University of Science and Technology of China. His research interests include computer graphics and non-image optics. In addition, he is also interested in autonomous driving and 3D AIGC.
[Uncaptioned image] Yuou Sun is currently working toward a PhD degree in the School of Mathematical Sciences, University of Science and Technology of China. His research interests include computer graphics, numerical optimization and non-image optics.
[Uncaptioned image] Bailin Deng is a senior lecturer in the School of Computer Science and Informatics at Cardiff University. He received the BEng degree in computer software (2005) and the MSc degree in computer science (2008) from Tsinghua University (China), and the PhD degree in technical mathematics (2011) from Vienna University of Technology (Austria). His research interests include geometry processing, numerical optimization, computational design, and digital fabrication. He is a member of the IEEE.
[Uncaptioned image] Juyong Zhang is a professor in the School of Mathematical Sciences at University of Science and Technology of China. He received the BS degree from the University of Science and Technology of China in 2006, and the Ph.D degree from Nanyang Technological University, Singapore. He mainly conducts research at the intersection of Vision, Graphics, and AI with a special focus on capturing, modeling and synthesizing objects, humans and large-scale scenes. He is an associate editor of IEEE Transactions on Mobile Computing and IEEE Computer Graphics and Applications.